OpenStudy (anonymous):

What is the exact value of the square root of the quantity 75 over 36.? A:5 the square root of 3 all over 6. B:25 the square root of 3 all over 9. C:5 the square root 3 all over 9. D:9 the square root of 3 all over 4.

OpenStudy (anonymous):

\[\sqrt{\frac{75}{36}} = \frac{\sqrt{3 \times 5 \times 5}}{\sqrt{6 \times 6}} = ???\]